# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

### v1

@@ 2 Upworthy Data and Experiments
We now discuss data sourced from Upworthy, a U.S. media publisher known for its innovative use of A/B test in online media publishing. In a bid to increase clicks to articles, Upworthy conducted randomized experiments with each article published, exploring different combinations of headlines and images to determine which elements most effectively engaged viewers. The archival record from January 24, 2013, to April 30, 2015, detailed in (matias2021upworthy), demonstrates how strategic packaging of headlines, subheadings, and images played a crucial role in Upworthy’s growth strategy. Such A/B testing practices are also widely used in the media and publishing industry, technology industry, and political campaigns. A/B testing allows firms to assess and refine the impact of various editorial and content choices (e.g., headlines or text of an article/email) systematically and automatically.
@@ 1 Why Yoko Ono Is The Only Thing Standing Between New York And Catastrophic Gas Fracking 2,734 34
During the experiment period, Upworthy’s editorial team created several versions of headlines and/or images for each article (internally called “package” by Upworthy). A package is defined as one treatment or arm for an article and consists of a headline, image, or a combination of both. Editors would first choose several of what they believed were the most promising packages to be tested. Then, they A/B tested these packages or treatments to identify which one resonated the best with their audience. Specifically, Upworthy’s A/B testing system recorded how many impressions and clicks each package received. Key metrics from these tests were shown via a dashboard, which guided decisions on whether to adjust the content further or finalize a particular version. Table 1 shows two examples of A/B tests, with the relevant columns. The full details of all the A/B tests and their results are available from the Upworthy Research Archive.
@@
Among these tests, some were conducted to test different images. However, the dataset does not allow us to trace back the actual images used in the tests. Therefore, we focus on the headline tests and filter out all the image tests. After this filtering, there are 17,681 tests comparing headlines, totaling 77,245 packages. On average, each article was tested with 4.37 headlines. Note that after filtering, each package or treatment simply refers to a unique headline; and therefore, we use the phrases package and headline interchangeably in the rest of the paper. The distribution of the number of packages/headlines for each article is shown in Table 3. After filtering, the dataset includes 277,338,713 impressions and 3,741,517 clicks; see the middle panel of Table 2. We also report the total number of impressions for the entire Upworthy website, as recorded in their data from Google Analytics. It is noteworthy that the impressions attributable to the tests constitute only 22.9% of the total traffic. This relatively small proportion of traffic can be attributed to Upworthy’s policy of conducting no more than one test per webpage to avoid interference effects.
@@
To examine whether these A/B tests were informative, or not, we generate pairs of headlines and conduct pairwise t-tests on the CTRs of the headlines. For example, if four headlines were tested for one article, we generate $4\times 3/2=6$ pairs of headlines, and test whether each pair is significantly different from each other; [^5] It is important to note that the issue of multiple testing may arise when conducting a large number of pairwise comparisons, as it increases the likelihood of Type I errors, or false positives. In our dataset, tests involving 10 pairs or less (equivalent to no more than 5 packages) account for over 80% of all tests. The risk of multiple testing is relatively low when conducting fewer than 10 tests, and conventional significance levels, e.g., $\alpha=0.05$ are generally adequate to account for this risk. To ensure the robustness of our findings, we conducted an additional analysis by excluding tests with more than 5 packages. The comparisons among pure-LLM-based methods remained consistent because all methods were conducted on the same dataset. It is important to note that a higher risk of multiple testing results in a larger number of actually insignificant pairs within our significant subset, which may only decrease the prediction accuracy. In other words, the results we present using the significant pairs may underestimate the prediction accuracy. Ultimately, our objective is to develop a more effective decision rule for assigning experimental traffic, and in Section 4, we utilize all available data without concern for multiple testing. see Table 4 for examples of such tests for the two A/B tests highlighted earlier in Table 1. After generating pairs, we obtain 140,655 comparable headline pairs across all tests, of which 39,158 pairs significantly differ in CTRs at a 0.05 level, and 23,682 at a 0.01 level. We do not observe a fundamental difference when conducting our experiments with 0.05 or 0.01 significant level pairs; therefore, for simplicity, we primarily use 0.05 level data in the main analysis in Section 3. In summary, we find that only 28% of the tests were significant at the 0.05 level, and a mere 17% are significant at the 0.01 level. Overall, these results suggest that most pairs of headlines tend to perform similarly, and the task of learning which headline is catchier is non-trivial.
@@ 3 Pure-LLM-Based Methods
In this section, we analyze 39,158 headline pairs, representing 28.05% of the total, at a 0.05 significance level. We focus on these significant pairs for three reasons: First, insignificant pairs suggest that decisions in final headline selection do not substantially affect users’ click behavior, implying that accuracy in predicting outcomes for “equally good” samples is not critically important from a business perspective. Second, to avoid information overload in our survey study, we limit our focus to significant pairs for a fair comparison between LLMs and humans. Third, the primary objective here is to evaluate various LLM-based classification methods to determine the most effective approach for the algorithm design in Section 4. As such, relative comparison across methods is more important as long as we use the same dataset for all methods. Furthermore, we note that the methods discussed in this section can be readily adapted from binary to multi-class classification tasks. Indeed, in Section 4, we incorporate practical elements such as multiple headlines and insignificant pairs to ensure our experimental setup closely mirrors real-world conditions and that our conclusions are directly applicable to business scenarios.
@@ 3.1 Prompt-based Approaches
In this section, we consider prompt-based approaches to evaluate which headline among a pair of headlines is catchier. These approaches simply use GPT prompts to learn which headline is catchier, and there is no further modeling or explicit fine-tuning involved. Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far .
We consider two prompt-based approaches: (1) Zero-shot prompting and (2) In-context learning. We discuss both in detail below.
@@ 3.1.1 Zero-shot Prompting
Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly trained on specific examples. In this study, we evaluate the performance of headline pair comparison using zero-shot prompting based on OpenAI’s GPT models. We do this through OpenAI’s API in Python, which offers various parameters to customize and control the model’s behavior and responses. The key parameters include specifying which GPT model to use (such as GPT-3 or GPT-4) and the system parameters, which guide the type of responses generated. The user’s input or query is contained in the user parameter, while the assistant parameter holds the model’s previous responses to maintain context in an ongoing conversation. Additionally, we set the default temperature parameter of 1.666The temperature parameter scales and controls the randomness of the model’s responses. GPT model generates responses token by token. Tokens are common sequences of characters found in a set of text. A helpful rule of thumb is that one token generally corresponds to about four characters of text in common English, translating to roughly three-quarters of a word (so 100 tokens $\approx$ 75 words), with each token sampled from a token distribution provided by the model. The temperature parameter determines whether this distribution is flattened or sharpened. A larger value results in a more diverse and creative output, while a lower temperature makes the output more deterministic and focused. For simplicity, we use the default setting with temperature as 1.0. We also tested a lower temperature of 0.1 to generate more deterministic outcomes, but we did not observe any significant difference in accuracy/performance.
@@
In our setting, we design the zero-shot prompting to ask GPT which headline is catchier, as shown in Figure 1. We test this prompt using three OpenAI large language models, [^7] As recommended by OpenAI, see this page, accessed on May 24, 2024, one can default to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o. OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence., each differing in parameter size: (1) GPT-3.5 (GPT-3.5-turbo-0125), which uses around 175 billion parameters, (2) GPT-4 (GPT-4-turbo-2024-04-09) which uses approximately 1.7 trillion parameters, and (3) GPT-4o (GPT-4o-2024-05-13), which uses an even larger number of parameters, enhancing its capabilities. The goal here is to investigate whether improvements in model architecture and parameter size can improve the accuracy of our predictions via standard prompts.
@@ 3.1.2 In-context Learning
In-context learning is a technique where a model is provided with demonstrations within the input prompt to help it perform a task. Demonstrations are sample inputs and outputs that illustrate the desired behavior or response. This method does not change the underlying parameters of the GPT models; rather, it leverages the model’s ability to recognize patterns and apply them to new, similar tasks by conditioning on input-output examples. Previous studies have shown that in-context learning can significantly boost a model’s accuracy and relevance in responses compared to zero-shot prompting. (SangMichaelXieandSewonMin_2022) explains this using a Bayesian inference framework, suggesting that in-context learning involves inferring latent concepts from pre-training data, utilizing all components of the prompt (inputs, outputs, formatting) to locate these concepts, even when training examples have random outputs. (xie2021explanation) further elucidate that in-context learning can emerge from models pre-trained on documents with long-range coherence, requiring the inference of document-level concepts to generate coherent text. They provide empirical evidence and theoretical proofs showing that both Transformers and LSTMs can exhibit in-context learning on synthetic datasets. [^8] More generally, beyond LLMs, (garg2022can) investigate the capacity of Transformers to in-context learn various function classes, demonstrating that these models can be trained from scratch to learn linear, sparse linear, two-layer neural networks, and decision trees, even under distribution shifts between training data and inference-time prompts.
We design the in-context learning prompt shown in Figure 2. Our prompts include demonstrations of catchier headlines and ask the model to determine which headline is catchier for a subsequent pair of headlines. We use the headline with the higher CTR from the test, as indicated in the “Winner” column of Table 4, as the correct answer in our demonstrations. A few demonstrations are typically sufficient to guide the model. Improvement tends to have diminishing returns quickly; adding more demonstrations does not significantly enhance performance after a certain point (min2022rethinking). In our numerical experiments, we use two or five demonstrations to investigate the impact of the number of examples on performance. [^9] We did run experiments with a larger number of demonstrations, but did not find any significant improvements from doing so; and hence do not report them here.
@@ 3.1.3 Performance Analysis of Prompt-based Approaches
Before presenting the results, we make note of two points. First, we do not find any significant difference between GPT-4 and GPT-4o; therefore, we only report the results from GPT-3.5 and GPT-4 for simplicity. Second, the GPT models can produce different answers even when given the exact same prompt. To mitigate this variability, we randomly select 1,000 significant pairs and conduct all our experiments on this same test set. Similarly, all the demonstrations are also randomly sampled and fixed for all experiments to reduce the variance of prompt outputs.
The results of all the prompt-based experiments are shown in Table 5. A more detailed pairwise comparison and t-test analysis of all the different experiments are available in Web Appendix A.1. The second column of 5, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting. Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with flipped outcome labels. The third column, “Flipped Label” indicates whether the labels were flipped or not.
@@
We find that prompt-based approaches using GPT-3.5, except for the zero-shot prompting, are not significantly better than random predictions (accuracy rates hover around 50%). Neither the number of demonstrations used nor the label type have any significant impact on the accuracy of the GPT-3.5 predictions. Interestingly, we find zero-shot prompting using GPT-3.5 is significantly better than the random prediction, although by only 4.20%, and marginally outperforms the in-context learning. In comparison, experiments with GPT-4 have accuracy rates around 60-65%, which is a significant improvement in performance compared to GPT-3.5. For GPT-4, in-context learning significantly improves performance over zero-shot prompting in GPT-4, though increasing the number of demonstrations from two to five does not result in a significant difference in performance. There is also no significant difference between using correct labels and incorrect labels in the demonstrations for in-context learning. This confirms that in-context learning helps GPT understand the task better through the input-output examples, but it fails to leverage the information in mapping from inputs to outputs effectively.
Finally, we also report the monetary cost and total run-time of the prompt-based experiments. The cost is \$1.13 for GPT-3.5 with 1,000 pairs and \$20.11 for GPT-4, which is approximately 20 times the cost of GPT-3.5. The total run-time is 121.5 minutes for the entire experiment, which involved testing a total of 10,000 prompts. [^10] Note that the run-time may depend on the API usage tiers, which provide different rate limits.
In summary, we find that prompt-based methods are relatively slow (compared to other methods introduced later in this section), have low accuracy (no more than 65%), high randomness (different answer outputs even under the same prompt inputs), and high monetary cost. All of these drawbacks make them unsuitable for real-time business applications, i.e., these methods cannot replace experimentation for content selection in digital platforms.
@@ 3.2 Classification Models with LLM-based Text Embeddings
In this section, we consider another LLM-based method for this prediction task. Specifically, we leverage the text embeddings from OpenAI and utilize classification models, including Logistic Regression (Logit) and Multilayer Perceptron (MLP), to predict which of the headlines in a pair is likely to generate a higher CTR. The Logit model is one of the simplest binary classification models and predicts the probability that a given input belongs to a particular class by applying the logistic function to a linear combination of the input variables. On the other hand, MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data. We refer interested readers to Chapter 5 in (zhang2021dive) for more details.
3.2.1 Text Embedding Technique
Text embedding is an advanced technique in natural language processing (NLP) that transforms larger chunks of text, such as sentences, paragraphs, or documents, into numerical vectors. Such vector forms a representation of the semantic information of the input text, and can be further utilized by downstream machine learning models.
Models like BERT (devlin2018bert), Sentence-BERT (reimers2019sentence), and OpenAI’s embedding models (openai-embedding) (text-embedding-3-large for example) are viewed as state-of-the-art embedding models at the date of their release. Essentially, text embedding is a numerical representation of text that captures the rich information in the text. These embedding vectors are used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which determines the emotional tone of a piece of text.
It is important to distinguish between embedding models and standard LLMs used for prompts, e.g., GPT. While they are based on similar underlying Transformer architectures and extensive pretraining on large text corpora, they are optimized for different tasks. Models like GPT are optimized for generating coherent and contextually relevant text based on input prompts, while embedding models focus on producing high-quality embeddings for downstream tasks. For a more detailed survey of the most recent text embedding models (patil2023survey).
In the remainder of this section, we leverage the embedding model from OpenAI to convert headline texts into numerical embedding vectors, which are then used to train the downstream binary classification models to identify which headline is better.

### v2

@@ 1 Why Yoko Ono Is The Only Thing Standing Between New York And Catastrophic Gas Fracking 2,734 34
We now discuss the data sourced from Upworthy, a U.S. media publisher known for its extensive use of A/B tests in digital publishing. Upworthy conducted randomized experiments with each article published, exploring different combinations of headlines and images to determine which elements most effectively led to higher clicks. The archival record from January 24, 2013, to April 30, 2015, detailed in (matias2021upworthy), demonstrates how the packaging of headlines and images played a crucial role in Upworthy’s growth strategy. Upworthy’s editorial team created several versions of headlines and/or images for each article (internally called “package” by Upworthy). A package is defined as one treatment or arm for an article and consists of a headline, image, or a combination of both. Editors would first choose several of what they believed were the most promising packages to be tested. Then, they A/B tested these packages or treatments to identify which one resonated the best with their audience. During the A/B test, users only saw the headline (and an accompanying image in some cases), but not the article itself. Upon clicking on the headline, they were taken to the article. As such, the content of the article itself did not have any effect on users’ clicking behavior. Upworthy’s A/B test system recorded how many impressions and clicks each package received. Table 1 shows two examples of A/B tests with the relevant columns. The full details of all the A/B tests and their results are available from the Upworthy Research Archive.
@@
Among these tests, some were conducted to test different images. However, the dataset does not allow us to trace back the actual images used in the tests. Therefore, we focus on the headline tests and filter out all the image tests. After this filtering, there are 17,681 tests, totaling 77,245 packages. On average, each article was tested with 4.37 headlines. Note that after filtering, each package or treatment simply refers to a unique headline; and therefore, we use the phrases package and headline interchangeably in the rest of the paper. The distribution of the number of packages/headlines for each article is shown in Table 3. After filtering, the dataset includes 277,338,713 impressions and 3,741,517 clicks; see the middle panel of Table 2. We also report the total number of impressions for the entire Upworthy website, as recorded in their data from Google Analytics. We see that the impressions attributable to the tests constitute only 22.9% of the total traffic.
@@
To examine whether these A/B tests were informative, or not, we generate pairs of headlines and conduct pairwise t-tests on the CTRs of the headlines. For example, if four headlines were tested for one article, we generate $4\times 3/2=6$ pairs of headlines, and test whether each pair is significantly different from each other; [^3] The issue of multiple testing may arise when conducting a large number of pairwise comparisons, as it increases the likelihood of Type I errors, or false positives. In our dataset, tests involving 10 pairs or less (equivalent to no more than 5 packages) account for over 80% of all tests. The risk of multiple testing is relatively low when conducting fewer than 10 tests, and conventional significance levels, e.g., $\alpha=0.05$ are generally adequate to account for this risk. Ultimately, our objective is to develop a more effective decision rule for assigning experimental traffic, and in §$\S$4, we utilize all available data without concern for multiple testing. see Table 4 for examples of such tests for the two A/B tests highlighted earlier in Table 1. After generating pairs, we obtain 140,655 comparable headline pairs across all tests, of which 39,158 pairs significantly differ in CTRs at a 0.05 level and 23,682 at a 0.01 level. In summary, we find that only 28% of the tests were significant at the 0.05 level, and a mere 17% are significant at the 0.01 level. Overall, these results suggest that most pairs of headlines tend to perform similarly, and the task of learning which headline is catchier is non-trivial. To further illustrate the difficulty of this task, we conduct a survey, where we give human users pairs of headlines from the set of headline pairs that are significantly different from each other and ask them to identify the catchier headline. Even within this set of headline pairs, we find that the respondents’ accuracy in the survey was not significantly better than random guessing. This further highlights the fact that the task of identifying engaging content is inherently challenging, even for humans. See Web Appendix §$\S$§A for more details of the survey and its results.
@@ 3 Pure-LLM-Based Methods
In this section, we consider the 39,158 headline pairs that significantly differ from each other at the 0.05 significance level. We focus on these significant pairs for three reasons: First, insignificant pairs suggest that decisions in final headline selection do not substantially affect users’ click behavior, implying that accuracy in predicting outcomes for “equally good” samples is not critically important from a business perspective. Second, the primary objective here is to evaluate various LLM-based classification methods to determine the most effective approach for the algorithm design in §$\S$4. As such, relative comparison across methods is more important as long as we use the same dataset for all methods. Third, as demonstrated in Web Appendix §$\S$§D.2.1, our robustness checks confirm that including insignificant pairs in the training phase does not affect the overall performance of the methods considered. Furthermore, we note that all methods discussed in this section can be readily adapted from binary to multi-class classification tasks. Indeed, in §$\S$4 where we introduce our LOLA approach, we incorporate practical elements such as multiple headlines and insignificant pairs to ensure our experimental setup closely mirrors real-world conditions and that our conclusions are directly applicable to business scenarios.
@@ 3.1 Prompt-based Approaches
Prompt-based approaches use GPT prompts to learn which headline is catchier, and there is no modeling or explicit fine-tuning involved. Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far and have been used to examine if and when LLMs can simulate consumer behavior in market research studies (gui2023challenge; brand2023using; li2024frontiers). We consider two prompt-based approaches: (1) Zero-shot prompting and (2) In-context learning.
@@ 3.1.1 Zero-shot Prompting
Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly trained on specific examples. We evaluate the performance of headline pair comparison using zero-shot prompting based on OpenAI’s GPT models. We do this through OpenAI’s API in Python, which offers various parameters to customize and control the model’s behavior and responses. The key parameters include specifying which GPT model to use (such as GPT-3 or GPT-4) and the system parameters, which guide the type of responses generated. The user’s input or query is contained in the user parameter, while the assistant parameter holds the model’s previous responses to maintain context in an ongoing conversation. Additionally, we set the default temperature parameter of 1.444The temperature parameter scales and controls the randomness of the model’s responses. A larger value results in a more diverse and creative output, while a lower temperature makes the output more deterministic and focused. For simplicity, we use the default setting with temperature as 1.0. We also tested a lower temperature of 0.1 to generate more deterministic outcomes, but we did not observe any significant difference in accuracy.
@@
We design the zero-shot prompting to ask GPT which headline is catchier, as shown in Figure 1. We test this prompt using three OpenAI large language models, [^5] As recommended by OpenAI, see this page, accessed on May 24, 2024, one can default to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o. OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence. each differing in parameter size: (1) GPT-3.5 (GPT-3.5-turbo-0125), which uses around 175 billion parameters, (2) GPT-4 (GPT-4-turbo-2024-04-09) which uses approximately 1.7 trillion parameters, and (3) GPT-4o (GPT-4o-2024-05-13), which uses an even larger number of parameters. The goal here is to investigate whether improvements in model architecture and parameter size can improve the accuracy of our predictions via standard prompts.
@@ 3.1.2 In-context Learning
In-context learning is a technique where a model is provided with demonstrations within the input prompt to help it perform a task. Demonstrations are sample inputs and outputs that illustrate the desired behavior or response. This method does not change the underlying parameters of the GPT models; rather, it leverages the model’s ability to recognize patterns and apply them to new, similar tasks by conditioning on input-output examples. Previous studies have shown that in-context learning can significantly boost a model’s accuracy and relevance in responses compared to zero-shot prompting. (SangMichaelXieandSewonMin_2022) explains this using a Bayesian inference framework, suggesting that in-context learning involves inferring latent concepts from pre-training data, utilizing all components of the prompt (inputs, outputs, formatting) to locate these concepts, even when training examples have random outputs. (xie2021explanation) further elucidate that in-context learning can emerge from models pre-trained on documents with long-range coherence, requiring the inference of document-level concepts to generate coherent text. They provide empirical evidence and theoretical proofs showing that both Transformers and LSTMs can exhibit in-context learning on synthetic datasets.
We design in-context learning prompts as shown in Figure 2. Our prompts include demonstrations of catchier headlines and ask the model to determine which headline is catchier for a subsequent pair of headlines. We use the headline with the higher CTR from the test, as indicated in the “Winner” column of Table 4, as the correct answer in our demonstrations. A few demonstrations are typically sufficient to guide the model. Improvement tends to have diminishing returns quickly, i.e., adding more demonstrations does not significantly enhance performance after a certain point (min2022rethinking). In our numerical experiments, we use two or five demonstrations to investigate the impact of the number of examples on performance. [^6] We did run experiments with a larger number of demonstrations, but did not find any significant improvements from doing so; and hence do not report them here.
@@ 3.1.3 Performance Analysis of Prompt-based Approaches
Before presenting the results, we make note of two points. First, we do not find any significant difference between GPT-4 and GPT-4o; therefore, we only report the results from GPT-3.5 and GPT-4 for simplicity. Second, the GPT models can produce different answers even when given the exact same prompt. To mitigate this variability, we randomly select 1,000 significant pairs of headlines and conduct all our experiments on this same test set. Similarly, all the demonstrations are also randomly sampled and fixed for all experiments to reduce the variance of prompt outputs.
The results of all the prompt-based experiments are shown in Table 5. A more detailed pairwise comparison and t-test analysis of all the different experiments are available in Web Appendix §$\S$§B.1. The second column of Table 5, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting. Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with flipped outcome labels. The third column, “Flipped Label” indicates whether the labels were flipped or not.
@@
We find that prompt-based approaches using GPT-3.5, except for the zero-shot prompting, are not significantly better than random predictions (accuracy rates hover around 50%). Neither the number of demonstrations used nor the label type have any significant impact on the accuracy of the GPT-3.5 predictions. Interestingly, zero-shot prompting using GPT-3.5 is slightly better than random prediction and also marginally outperforms in-context learning. In comparison, experiments with GPT-4 have accuracy rates around 60-65%, which is a significant improvement in performance compared to GPT-3.5. For GPT-4, in-context learning significantly improves performance over zero-shot prompting in GPT-4, though increasing the number of demonstrations from two to five does not result in a significant difference in performance. There is also no significant difference between using correct labels and incorrect labels in the demonstrations for in-context learning. This confirms that in-context learning helps GPT understand the task better through the input-output examples, but it fails to leverage the information in mapping from inputs to outputs effectively.
In summary, we find that prompt-based methods have low accuracy (no more than 65%), high randomness (different answer outputs even under the same prompt inputs), slow running time (compared to other methods introduced later in this section), and high monetary cost (see Web Appendix §$\S$§E for a detailed discussion of costs and running time). All of these drawbacks make them unsuitable for real-time business applications, i.e., these methods cannot replace experimentation for content selection in digital platforms.
@@ 3.2 Classification Models with LLM-based Text Embeddings
We now consider another LLM-based method for the prediction task. Specifically, we leverage the text embeddings from OpenAI and utilize classification models, including Logistic Regression (Logit) and Multilayer Perceptron (MLP), to predict which of the headlines in a pair is likely to generate a higher CTR.
Text embedding techniques transform large chunks of text, such as sentences, paragraphs, or documents, into numerical vectors. These embedding vectors can then be used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which is used to determine the emotional tone of a piece of text. [^7] Embedding models and standard prompt-based LLMs used share some common features, e.g., the underlying Transformer architectures and extensive pretraining on large text corpora. However, they are optimized for different tasks. Models like GPT are optimized for generating coherent and contextually relevant text based on input prompts while embedding models focus on producing high-quality embeddings for downstream tasks. See patil2023survey for a more detailed survey of the most recent text embedding models and their applications.
An interesting question is why text embeddings obtained from LLM models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for downstream classification tasks. (saunshi2020mathematical) provide an explanation by demonstrating that classification tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text. Thus, the embeddings, which can be roughly viewed as the next token probability generated from a next-word prediction task, inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).